OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2

OpenStudy (agreene):

your slope is the change in y over the change in x or \[\Delta y/\Delta x\] so, with these two points, we plug in and solve for x -3/2 = (-3-6)/(10-x) -3/2 = -9/(10-x) multiply both sides by the denominator (10-x)*-3/2 = -9 divide by -3/2 10-x = -9/(-3/2) 10-x=6 subtract over your 10 -x= -4 multiply both sides by -1 x=4 :)